OpenStudy (anonymous):

F(x)= 2x^3+x^2 G(x)= 7x-2 Find: f(3)/g

jhonyy9 (jhonyy9):

f(3)=2*27 +9=54+9=63 f(3)/g = 63/(7x-2)